A spherical-shaped capacitor has a charge equivalent to 4.45 nanocoulombs when connected to a battery that has a voltage of 220 V. Suppose the two shells of the spherical capacitor are 5.34 centimeters apart, what are the capacitance and the radius of the inner sphere?

1
Expert's answer
2022-03-15T10:34:37-0400

The capacitance is given by

"C=\frac{q}{V}=\frac{4.45*10^{-9}}{220}=2.02*10^{-11}\:\rm F"

On the other hand

"C=4\pi\epsilon_0(R_2*R_1)/(R_2-R_1)""R_2=R_1+0.0534"

So

"2.02*10^{-11}=3.14*8.85*10^{-12}\frac{(R_1+0.0534)R_1}{0.0534}"

"(R_1+0.0534)R_1=0.0389"

"R_1=0.172\:\rm m"

"R_2=0.226\:\rm m"